OpenStudy (anonymous):

this is a bunch of questions that i cannot figure out i will give medals to whoever can answer them and number them... please oh please oh please :) What is the value of the coefficient "b" when the quadratic equation y = (3x - 5)(2x + 3) is written in standard form? Identify the vertex for the graph of y = -3x^2 - 6x + 4. Determine whether the graph of y = -3x^2 + 2x - 8 opens up or down and whether it has a maximum or minimum point. Determine whether the graph of y = x^2 - 6x + 3 has a maximum or minimum point, then find the maximum or minimum value.

OpenStudy (nenadmatematika):

first question: when you multiply you get 6x^2-x-15 so b=-1

OpenStudy (nenadmatematika):

second question: one way to find the vertex is to find the first derivative of that function and set it to be zero, so: -6x-6=0 x=-1 plug back in x=-1 into your function and you get: y=7 so your vertex is (-1,7)

OpenStudy (nenadmatematika):

third question: since coefficient a=-3 your function is concave, so it has maximum

OpenStudy (nenadmatematika):

fourth question: the same pattern: a>0 so the function is convex, it has a minimum point and that is the vertex: so 2x-6=0 x=3 plug back in and y=-6 so the minimum is (3,-6)